OpenStudy (anonymous):

What is the area of a square with a perimeter of 14?

OpenStudy (anonymous):

Perimeter = 4*side length. 14/4 = side Area = side^2 Area = (14/4)^2=12.25 or 49/4

OpenStudy (anonymous):

both of those are a choice so are they both right?

OpenStudy (iwantanswers):

rhombus

OpenStudy (anonymous):

who are you iwantanswers

OpenStudy (anonymous):

One is in fractional form and the other is in decimal. Some teachers want the decimal answer. Others like the fractional form since they want proof of work.

OpenStudy (anonymous):

Plug 49 and divide it by 4 in a calculator and see what you get :)

OpenStudy (iwantanswers):

RHOMBUS!